# Self-embedding similitudes of Bedford-McMullen carpets with dependent ratios

### v1

@@ Abstract.
We prove that any non-degenerate Bedford-McMullen carpet does not allow oblique self-embedding similitudes; that is, if $f$ is a similitude sending the carpet into itself, then the image of the $x$-axis under $f$ must be parallel to one of the principal axes. We also establish a logarithmic commensurability result on the contraction ratios of such embeddings. This completes a previous study of Algom and Hochman [Ergod. Th. & Dynam. Sys. 39 (2019), 577–603] on Bedford-McMullen carpets generated by multiplicatively independent exponents, together with a new proof on their non-obliqueness statement.
@@
Bedford-McMullen carpets, self-embedding, obliqueness, generalized Sierpiński carpets, deleted-digit sets.
@@ 2.1. Bedford-McMullen carpets
Note for every $y\in\pi_{2}(K)$, say $y=\sum_{k=1}^{\infty}y_{k}m^{-k}$ where $y_{k}\in J$,
@@
The following lemma records some basic facts about horizontal slices.
@@
Let $K=K(n,m,\Lambda)$ be a Bedford-McMullen carpet. Then for any $y=\sum_{k=1}^{\infty}y_{k}m^{-k}$ with $\{y_{k}\}_{k=1}^{\infty}\subset J$,
@@ 2.2. Deleted-digit sets
Let $K=K(n,m,\Lambda)$ be a Bedford-McMullen carpet with $12\leq N\leq n-$ 1 and let $y=\sum_{k=1}^{\infty}y_{k}m^{-k}$ with $\{y_{k}\}_{k=1}^{\infty}\subset J$. Let $\alpha:=\frac{\log N}{\log n}$. If $0<\mathcal{H}^{\alpha}(K^{y})<\infty$ and there is a contraction $h\in\mathcal{S}(\mathbb{R})$ sending $K^{y}$ into itself, then the contraction ratio of $h$ is a rational power of $n$.
@@
We first prove that #$\#I_{y_{k}}=N$ for all large $k$. Suppose on the contrary that #$\#I_{y_{k}}<N$ for infinitely many $k$. Then for any $p\geq 1$, when $\widetilde{M}$ is large enough, we have
@@
Let $K,N,y,h$ be as in Proposition 2.6. Then there are left isolated points $a,a^{\prime}$ and right isolated points $b,b^{\prime}$ of $K^{y}$ such that $h(a)=a^{\prime}$ and $h(b)=b^{\prime}$. Moreover, we can guarantee that $\pi_{2}(R_{t})=\pi_{2}(R^{\prime}_{t})$ for all large $t$, where $R_{t},R^{\prime}_{t}$ are the level-$t$ rectangles such that $(a,y)\in\varphi_{R_{t}}(K)$ and $(b,y)\in\varphi_{R^{\prime}_{t}}(K)$, respectively.
@@
For the second one, note that if $y\notin\bigcup_{t=1}^{\infty}\frac{\mathbb{Z}}{m^{t}}\cap[0,1]$, then $R_{t}$ and $R^{\prime}_{t}$ are unique for all $t$. Moreover, $y$ belongs to the interior of both of $\pi_{2}(R_{t})$ and $\pi_{2}(R^{\prime}_{t})$, which implies that $\pi_{2}(R_{t})=\pi_{2}(R^{\prime}_{t})$. If $y\in\bigcup_{t=1}^{\infty}\frac{\mathbb{Z}}{m^{t}}\cap[0,1]$, it is possible that $\pi_{2}(R_{t})=\pi_{2}(R^{\prime}_{t})+m^{-t}$ or vice versa (note that $\pi_{2}(R_{t}),\pi_{2}(R^{\prime}_{t})$ are both $n$-adic intervals of length $m^{-t}$). However, for such a point $y$, it is easy to see that $K^{y}$ is a finite union of scaled copies of $K^{0}$ and $K^{1}$ with disjoint interior. More precisely, $K^{y}=\bigcup_{s\in\Omega}(K^{0,s}\cup K^{1,s})$, where $K^{0,s}$ (resp. $K^{1,s}$) are scaled copies of $K^{0}$ (resp. $K^{1}$) contracting by a factor $m^{-t-1}$. By Baire’s theorem, some $K^{0,s}$ or $K^{1,s}$, say $K^{0,s}$, must contain an (relative) interior part of $h^{q}(B\cap K^{y})$ and hence of $h^{q}(B\cap K^{0,s})$. Then, picking those points so carefully that $a^{\prime},b^{\prime},(h^{q})^{-1}(a^{\prime}),(h^{q})^{-1}(b^{\prime})$ are all isolated points of $K^{0,s}$, we can guarantee that $\pi_{2}(R_{t})=\pi_{2}(R^{\prime}_{t})$ for all large $t$. This completes the proof. ∎
@@ 3. The independent case
Let us pick a sequence $\{R_{k}\}_{k=1}^{\infty}$, where each $R_{k}$ is a level-$k$ rectangle and $R_{k+1}\subset R_{k}$ for all $k\geq 1$. Write $a_{k}$ to be the left bottom vertex of $R_{k}$ and $Q_{k}$ to be any level-$k$ rectangle containing $f(a_{k})$. For convenience, we also let $\ell^{k},\ell_{k}$ be the bottom and left edges of $R_{k}$, respectively. So $\ell^{k}\cap\ell_{k}=\{a_{k}\}$. See Figure 2 for an illustration. Below we record two simple but key facts.
@@
    $=\sqrt{n^{2k}\cdot\lambda^{2}|\ell^{k}|^{2}\cdot\frac{1}{1+u^{2}}+m^{2k}\cdot\lambda^{2}|\ell^{k}|^{2}\cdot\frac{u}{1+u^{2}}}$
    $=\sqrt{n^{2k}\cdot\frac{\lambda^{2}}{n^{2k}}\cdot\frac{1}{1+u^{2}}+m^{2k}\cdot\frac{\lambda^{2}}{n^{2k}}\cdot\frac{u}{1+u^{2}}}$
@@
    $=\sqrt{n^{2k}\cdot\frac{\lambda^{2}}{m^{2k}}\cdot\frac{1}{1+u^{-2}}+m^{2k}\cdot\frac{\lambda^{2}}{m^{2k}}\cdot\frac{u^{-1}}{1+u^{-2}}}$
@@
Write $\widetilde{a}_{k}:=\varphi_{Q_{k}}^{-1}f(a_{k})$. Since $f(a_{k})\in Q_{k}$, $\widetilde{a}_{k}\in[0,1]^{2}$. Fix a large $k$ so that $\frac{m^{k}}{n^{k}}\cdot\max\{u^{-1},u\}<\lambda$ and $|\varphi_{Q_{k}}^{-1}f(\ell^{k})|\leq\frac{2\lambda}{\sqrt{1+u^{2}}}$ (recall (3.2)). Let us truncate the parallelogram $\varphi_{Q_{k}}^{-1}f(R_{k})$ as $\bigcup_{t\geq 0}\widetilde{E}_{t}$, where
@@
Since $\varphi_{Q_{k}}^{-1}(f(\ell_{k}))$ has slope $-\frac{m^{k}}{n^{k}u}$, we have for every $t$ that
(3.3) $|\pi_{2}(\widetilde{E}_{t})|\leq 2\Big{|}\frac{m^{k}}{n^{k}u}\Big{|}+|\varphi_{Q_{k}}^{-1}f(\ell^{k})|<2\lambda+\frac{2\lambda}{\sqrt{1+u^{2}}}\leq 4\lambda<\frac{1}{6m}.$
Since $\widetilde{a}_{k}\in[0,1]^{2}$, $\pi_{2}(\widetilde{E}_{0})\subset[-\frac{1}{6m},1+\frac{1}{6m}]$. Combining this with (3.3), when $k$ is picked large enough at the beginning, we can find integers $t_{0}\geq 0$ and $0\leq j_{0}\leq m$ such that $|\pi_{1}(\widetilde{E}_{t_{0}})|=1$ while
$\widetilde{E}_{t_{0}}\subset\{x\in\mathbb{R}^{2}:\tfrac{j_{0}}{m}+\tfrac{1}{3m}<\pi_{2}(x)<\tfrac{j_{0}+1}{m}-\tfrac{1}{3m}\}.$
Let $j^{\prime}_{0}:=j_{0}\pmod{m-1}$. Note that $\frac{j_{0}}{m-1}$ is either 00 or $1$.
On the other hand, $\varphi_{Q_{k}}^{-1}f(\ell_{k})$ is supported on a line of slope $-\frac{m^{k}}{n^{k}u}$, say $g(x)=-\frac{m^{k}}{n^{k}u}x+c$. Since #$\#I_{j^{\prime}_{0}}<n$, there is some $10\leq i_{0}\leq n-$ 1 such that $K$ does not intersect the open rectangle $(\frac{i_{0}}{n},\frac{i_{0}+1}{n})\times(\frac{j^{\prime}_{0}}{m},\frac{j^{\prime}_{0}+1}{m})=:U$. Write $e$ to be the point $(-t_{0},\lfloor\frac{j_{0}}{m-1}\rfloor)$ and $x_{0}:=-t_{0}+\frac{i_{0}+1/2}{n}$. Then the point $\xi_{0}:=(x_{0},g(x_{0}))$ lies on the segment $\varphi_{Q_{k}}^{-1}f(\ell_{k})$, and
$U+e$ $=\Big{(}\frac{i_{0}}{n}-t_{0},\frac{i_{0}+1}{n}-t_{0}\Big{)}\times\Big{(}\frac{j^{\prime}_{0}}{m}+\lfloor\frac{j_{0}}{m-1}\rfloor,\frac{j^{\prime}_{0}+1}{m}+\lfloor\frac{j_{0}}{m-1}\rfloor\Big{)}$
@@
So it is easy to check that the ball centered at $\xi_{0}:=(x_{0},g(x_{0}))$ and of radius $|\varphi_{Q_{k}}^{-1}f(\ell^{k})|\leq 2\lambda<\frac{2}{9mn}$ is contained in $U+e$. In particular,
@@
Recall that $\ell^{k}$ is the bottom edge of the level-$k$ rectangle $R_{k}$. Since $\xi_{0}+\varphi_{Q_{k}}^{-1}f(\ell^{k})$ is parallel to $\varphi_{Q_{k}}^{-1}f(\ell^{k})$ and is contained in $\varphi_{Q_{k}}^{-1}f(R_{k})$ (the original flat parallelogram), there is some $y_{0}$ such that
(3.5) $(\varphi_{Q_{k}}^{-1}f(\ell^{k})+\xi_{0})\cap\varphi_{Q_{k}}^{-1}f\varphi_{R_{k}}(K)=\varphi_{Q_{k}}^{-1}f\varphi_{R_{k}}(K^{y_{0}})\neq\varnothing,$
where the non-empty conclusion is because $K^{y}\neq\varnothing$ for all $y$ (as pointed out at the beginning). Thus
$\varnothing$ $\neq(\varphi_{Q_{k}}^{-1}f(\ell^{k})+\xi_{0})\cap\varphi_{Q_{k}}^{-1}f\varphi_{R_{k}}(K)$ (just (3.5))
@@
which leads to a contradiction. ∎

### v2

@@ Abstract.
We prove that any non-degenerate Bedford-McMullen carpet does not allow oblique self-embedding similitudes; that is, if $f$ is a similitude sending the carpet into itself, then the image of the $x$-axis under $f$ must be parallel to one of the principal axes. We also establish a logarithmic commensurability result on the contraction ratios of such embeddings. This completes a previous study by Algom and Hochman [Ergod. Th. & Dynam. Sys. 39 (2019), 577–603] on Bedford-McMullen carpets generated by multiplicatively independent exponents, together with a new proof of their non-obliqueness statement.
@@
Bedford-McMullen carpets, self-embedding, obliqueness, generalized Sierpiński carpets, logarithmic commensurability, deleted-digit sets.
@@ 2.1. Bedford-McMullen carpets
Note for every $y\in\pi_{2}(K)$ with a unique expansion $y=\sum_{k=1}^{\infty}y_{k}m^{-k}$, where $y_{k}\in J$,
@@
The following lemma records some basic facts about these horizontal slices.
@@
Let $K=K(n,m,\Lambda)$ be a Bedford-McMullen carpet. Then for any $y\in\pi_{2}(K)$ with a unique expansion $y=\sum_{k=1}^{\infty}y_{k}m^{-k}$, $\{y_{k}\}_{k=1}^{\infty}\subset J$,
@@ 2.2. Deleted-digit sets
Let $K=K(n,m,\Lambda)$ be a Bedford-McMullen carpet with $12\leq N\leq n-$ 1 and let $y\in\pi_{2}(K)$. Write $\alpha:=\frac{\log N}{\log n}$. If $0<\mathcal{H}^{\alpha}(K^{y})<\infty$ and there is a contraction $h\in\mathcal{S}(\mathbb{R})$ sending $K^{y}$ into itself, then the contraction ratio of $h$ is a rational power of $n$.
@@
Without loss of generality, assume that $h$ is orientation-preserving. If $y\in\pi_{2}(K)$ has more than one expansion with coefficients in $J$, $y\in\bigcup_{k\geq 1}\frac{\mathbb{Z}}{m^{k}}\cap[0,1]$. In this case, it is easy to see that $K^{y}$ is a finite union of scaled copies of $K^{0}$ and $K^{1}$, say $K^{y}=\bigcup_{s\in\Omega}K^{0,s}\cup\bigcup_{t\in\Gamma}K^{1,t}$, where $\Omega,\Gamma\subset\mathbb{Z}$ and
$K^{0,s}:=\frac{K^{0}+s}{n^{k}}\quad\text{and}\quad K^{1,s}:=\frac{K^{1}+t}{n^{k}}$
for some large integer $k$. Since $h(K^{y})\subset K^{y}$, by Baire’s theorem, some $K^{0,s}$ or $K^{1,s}$, say $K^{0,s}$, must contain an (relative) interior part of $h(K^{y})$ and hence of $h(K^{0,s})$. So we can find a scaled copy $E$ of $K^{0}$ contracting by a factor $n^{-q}$ for some large integer $q$ such that $h(E)\subset K^{0,s}$. Since $K^{0}=E(n,I_{0})$ is a deleted-digit set with positive Hausdorff dimension, by Lemma 2.5, the contraction ratio of $h$ is a rational power of $n$.
Now suppose $y$ has a unique expansion $y=\sum_{k=1}^{\infty}y_{k}m^{-k}$, $\{y_{k}\}\subset J$. We first prove that #$\#I_{y_{k}}=N$ for all large $k$. Suppose on the contrary that #$\#I_{y_{k}}<N$ for infinitely many $k$. Then for any $p\geq 1$, when $\widetilde{M}$ is large enough, we have
@@
Let $K,N,y,h$ be as in Proposition 2.6. Then there are left and right isolated points $a,b$ of $K^{y}$ such that $h(a),h(b)$ (or $h(b),h(a)$ if $h$ is orientation-reversing) are also left and right isolated points of $K^{y}$. Moreover, we can guarantee that $\pi_{2}(R_{t})=\pi_{2}(R^{\prime}_{t})$ for all large $t$, where $R_{t},R^{\prime}_{t}$ are two level-$t$ rectangles with $(a,y)\in\varphi_{R_{t}}(K)$ and $(b,y)\in\varphi_{R^{\prime}_{t}}(K)$, respectively.
@@
Suppose first that $y$ has a unique expansion with coefficients in $J$. For the first statement, it suffices to prove it when $h$ is orientation-preserving. In fact, if $h$ is orientation-reversing, then $h^{2}$ is orientation-preserving. If $a,h^{2}(a)$ are both left (resp. right) isolated points of $K^{y}$, then $h(a)$ must be a right (resp. left) isolated point and we are done.
@@
For the second one, since $y$ has a unique expansion, $R_{t}$ and $R^{\prime}_{t}$ are unique for all $t$. Moreover, $y$ belongs to the interior of both of $\pi_{2}(R_{t})$ and $\pi_{2}(R^{\prime}_{t})$, which implies that $\pi_{2}(R_{t})=\pi_{2}(R^{\prime}_{t})$.
Now, suppose $y$ has more than one expansion. Similarly as in the proof of Proposition 2.6, we can write $K^{y}=\bigcup_{s\in\Omega}K^{0,s}\cup\bigcup_{t\in\Gamma}K^{1,t}$, where $\Omega,\Gamma\subset\mathbb{Z}$ and
$K^{0,s}:=\frac{K^{0}+s}{n^{k}}\quad\text{and}\quad K^{1,s}:=\frac{K^{1}+t}{n^{k}}$
for some large integer $k$. By Baire’s theorem again, we may assume without loss of generality that some $K^{0,s}$ contains an interior part of $h(K^{y})$. In particular, there is a scaled copy of $K^{0,s}$, say $E=\frac{K^{0}+z}{n^{p}}\subset K^{0,s}$, such that $h(E)\subset K^{0,s}$. Note that
$h(E)\subset K^{0,s}\Longleftrightarrow h\Big{(}\frac{K^{0}+z}{n^{p}}\Big{)}\subset\frac{K^{0}+s}{n^{k}}.$
We can easily reformulate this as $g(K^{0})\subset K^{0}$ for some similitude $g$ with contraction ratio $\rho n^{k-p}$. Since 00 has a unique expansion, as in the first part of this proof, there are nondegenerate open intervals $U,V$ such that $g(U\cap K^{0})=V\cap K^{0}\neq\varnothing$ (recall (2.13)). In particular, $g$ maps all the one-sided isolated points of $K^{0}$ in $U$ to those of $K^{0}$ in $V$. Since the inclusion $g(K^{0})\subset K^{0}$ is just a reformulation of $h(E)\subset K^{0,s}$, $h$ maps all one-sided isolated points of $K^{0,s}$ in $U^{\prime}$ (for some $U^{\prime}$) to one-sided isolated points of $K^{0,s}$ in $V^{\prime}$ (for some $V^{\prime}$). It remains to prove that these points are also one-sided isolated points of $K^{y}$.
To get the desired statement, note first that we may pick $V^{\prime}$ properly so that
(2.14) $K^{0,s}\cap V^{\prime}=K^{y}\cap V^{\prime}.$
In fact, if $t\neq s$ for all $t\in\Gamma$, then we already have (2.14). If some $t$ equals $s$, then $K^{0,s}$ and $K^{1,t}=K^{1,s}$ may overlap. However, note that $\operatorname{dim_{H}}K^{0,s}\geq\operatorname{dim_{H}}K^{y}\geq\operatorname{dim_{H}}K^{1,s}$ (recall that $h(K^{y})\subset K^{0,s}$), we have either $K^{1,s}=K^{0,s}$ or these two deleted-digit sets have different patterns (in particular, $K^{0,s}\neq K^{1,s}$). So there is a small open interval $V^{\prime\prime}\subset V^{\prime}$ such that $(K^{0,s}\cap V^{\prime\prime})\cap K^{1,s}=\varnothing$ but $K^{0,s}\cap V^{\prime\prime}\neq\varnothing$. Replacing $V^{\prime}$ with $V^{\prime\prime}$ gives us (2.14). In particular, all left (resp. right ) isolated points of $K^{0,s}$ in $V^{\prime}$ are left (resp. right) isolated points of $K^{y}$.
Next, if a one-sided isolated point $a$ of $K^{0,s}$ in $U^{\prime}$ is not a one-sided isolated point of $K^{y}$, there is a sequence $\{x_{n}\}\subset K^{y}$ such that $x_{n}\to a$ from the left or from the right. Then $h(x_{n})\to h(a)$ from one direction. Since $h(x_{n})\in K^{y}$, this contradicts the fact that $h(a)$ is a one-sided isolated point of $K^{y}$ in $V^{\prime}$, which completes the proof. ∎
@@ 3. The independent case
Let us pick a sequence $\{R_{k}\}_{k=1}^{\infty}$, where each $R_{k}$ is a level-$k$ rectangle and $R_{k+1}\subset R_{k}$ for all $k\geq 1$. Write $a_{k}$ to be the left bottom vertex of $R_{k}$ and $Q_{k}$ to be any $n^{-k}\times m^{-k}$ grid rectangle containing $f(a_{k})$. For convenience, we also let $\ell^{k},\ell_{k}$ be the bottom and left edges of $R_{k}$, respectively. So $\ell^{k}\cap\ell_{k}=\{a_{k}\}$. See Figure 2 for an illustration. Below we record two simple but key facts. Here we abuse notation slightly by writing $\varphi_{Q_{k}}$ to be the natural affine map sending $[0,1]^{2}$ onto $Q_{k}$.
@@
    $=\sqrt{n^{2k}\cdot\lambda^{2}|\ell^{k}|^{2}\cdot\frac{1}{1+u^{2}}+m^{2k}\cdot\lambda^{2}|\ell^{k}|^{2}\cdot\frac{u^{2}}{1+u^{2}}}$
    $=\sqrt{n^{2k}\cdot\frac{\lambda^{2}}{n^{2k}}\cdot\frac{1}{1+u^{2}}+m^{2k}\cdot\frac{\lambda^{2}}{n^{2k}}\cdot\frac{u^{2}}{1+u^{2}}}$
@@
    $=\sqrt{n^{2k}\cdot\frac{\lambda^{2}}{m^{2k}}\cdot\frac{1}{1+u^{-2}}+m^{2k}\cdot\frac{\lambda^{2}}{m^{2k}}\cdot\frac{u^{-2}}{1+u^{-2}}}$
@@
Write $\widetilde{a}_{k}:=\varphi_{Q_{k}}^{-1}f(a_{k})$. Since $f(a_{k})\in Q_{k}$, $\widetilde{a}_{k}\in[0,1]^{2}$. Fix a large $k$ (will be specified later) so that $\frac{m^{k}}{n^{k}}\cdot\max\{u^{-1},u\}<\lambda$ and $|\varphi_{Q_{k}}^{-1}f(\ell^{k})|\leq\frac{2\lambda}{\sqrt{1+u^{2}}}$ (recall (3.2)). Let us truncate the parallelogram $\varphi_{Q_{k}}^{-1}f(R_{k})$ as $\bigcup_{t\in\mathbb{Z}}\widetilde{E}_{t}$, where
@@
Note that for every $t$,
$|\pi_{2}(\widetilde{E}_{t})|$ $\leq|\pi_{2}(\varphi_{Q_{k}}^{-1}f(R_{k}))|$
$\leq|\pi_{2}(\varphi_{Q_{k}}^{-1}f(\ell^{k}))|+|\pi_{2}(\varphi_{Q_{k}}^{-1}f(\ell_{k}))|$
(3.3) $=m^{k}\cdot\frac{\lambda}{n^{k}}\cdot\frac{u}{\sqrt{1+u^{2}}}+m^{k}\cdot\frac{\lambda}{m^{k}}\cdot\frac{u^{-1}}{\sqrt{1+u^{-2}}}<2\lambda.$
Let us prove Case 3 under the following Condition A first and then explain how to move on without its validity.
Condition A
There are integers $t_{0}\geq-1$ and $0\leq j_{0}\leq m$ and such that $|\pi_{1}(\widetilde{E}_{t_{0}})|=1$ while $\pi_{2}(\widetilde{E}_{t_{0}})\subset[\tfrac{j_{0}}{m},\tfrac{j_{0}+1}{m}]$.
Assume Condition A. Let $j^{\prime}_{0}:=j_{0}\pmod{m}$. Recall that $\varphi_{Q_{k}}^{-1}f(\ell_{k})$ is supported on a line of slope $-\frac{m^{k}}{n^{k}u}$, say $g(x)=-\frac{m^{k}}{n^{k}u}x+c$. Since #$\#I_{j^{\prime}_{0}}<n$, there is some $10\leq i_{0}\leq n-$ 1 such that $K$ does not intersect the open rectangle $(\frac{i_{0}}{n},\frac{i_{0}+1}{n})\times(\frac{j^{\prime}_{0}}{m},\frac{j^{\prime}_{0}+1}{m})=:U$. Write $e$ to be the point $(-t_{0},\lfloor\frac{j_{0}}{m}\rfloor)$ and $x_{0}:=-t_{0}+\frac{i_{0}+1/2}{n}$. Then the point $\xi_{0}:=(x_{0},g(x_{0}))$ lies on the segment $\varphi_{Q_{k}}^{-1}f(\ell_{k})$, and
$U+e$ $=\Big{(}\frac{i_{0}}{n}-t_{0},\frac{i_{0}+1}{n}-t_{0}\Big{)}\times\Big{(}\frac{j^{\prime}_{0}}{m}+\lfloor\frac{j_{0}}{m}\rfloor,\frac{j^{\prime}_{0}+1}{m}+\lfloor\frac{j_{0}}{m}\rfloor\Big{)}$
@@
Since $\pi_{2}(\widetilde{E}_{t_{0}})\subset[\tfrac{j_{0}}{m},\tfrac{j_{0}+1}{m}]$, it is not hard to check that
@@
Since $\xi_{0}+\varphi_{Q_{k}}^{-1}f(\ell^{k})$ is a translated copy of $\varphi_{Q_{k}}^{-1}f(\ell^{k})$ and is contained in $\varphi_{Q_{k}}^{-1}f(R_{k})$ (the original flat parallelogram), $f^{-1}\varphi_{Q_{k}}(\xi_{0})+\ell^{k}$ is a translated copy of $\ell^{k}$ and is contained in $R_{k}$. Since $\ell^{k}$ is the bottom edge of the level-$k$ rectangle $R_{k}$, $\ell^{k}\cap\varphi_{R_{k}}(K)=\varphi_{R_{k}}(K^{0})$. So there exists some $y_{0}\in[0,1]$ such that
$(f^{-1}\varphi_{Q_{k}}(\xi_{0})+\ell^{k})\cap\varphi_{R_{k}}(K)=\varphi_{R_{k}}(K^{y_{0}})\neq\varnothing,$
where the non-empty conclusion is because $K^{y}\neq\varnothing$ for all $y$ (as pointed out at the beginning of Case 3). This in turn implies that
(3.5) $(\xi_{0}+\varphi_{Q_{k}}^{-1}f(\ell^{k}))\cap\varphi_{Q_{k}}^{-1}f\varphi_{R_{k}}(K)=\varphi_{Q_{k}}^{-1}f\varphi_{R_{k}}(K^{y_{0}})\neq\varnothing.$
Thus
$\varnothing$ $\neq(\xi_{0}+\varphi_{Q_{k}}^{-1}f(\ell^{k}))\cap\varphi_{Q_{k}}^{-1}f\varphi_{R_{k}}(K)$ (just (3.5))
@@
which leads to a contradiction.
Now, suppose Condition A is not fulfilled. Set $j_{*}:=\min\{j\in\mathbb{Z}:\pi_{2}(\widetilde{a}_{k})<\frac{j}{m}\}\geq 1$. So the segment $[-2,-1]\times\{\frac{j_{*}}{m}\}$ should intersect $\widetilde{E}_{2}$: otherwise, Condition A holds for $\widetilde{E}_{2}$ and $j_{*}$ or $j_{*}-1$. We claim that
(3.6) $\min\pi_{2}(\widetilde{E}_{2}\cap(\{-2\}\times\mathbb{R}))<\tfrac{j_{*}}{m}<\max\pi_{2}(\widetilde{E}_{2}\cap(\{-1\}\times\mathbb{R})).$
Otherwise, either $\max\pi_{2}(\widetilde{E}_{2}\cap(\{-1\}\times\mathbb{R}))\leq\frac{j_{*}}{m}$ and thus Condition A holds for $\widetilde{E}_{1}$ and $j_{*}-1$; or $\min\pi_{2}(\widetilde{E}_{2}\cap(\{-2\}\times\mathbb{R}))\geq\frac{j_{*}}{m}$ and thus Condition A holds for $\widetilde{E}_{3}$ and $j_{*}$.
If #$\#I_{j^{\prime}}\geq 2$ for some $j^{\prime}$, then we can pick $R_{k}$ properly at the beginning such that
$R_{k}=\varphi_{i_{1},j^{\prime}}\varphi_{i_{2},j^{\prime}}\cdots\varphi_{i_{k},j^{\prime}}([0,1]^{2}),$
where $i_{1},\ldots,i_{k}\in I_{j^{\prime}}$. As a result, selecting $i^{\prime}_{k}\in I_{j^{\prime}}\setminus\{i_{k}\}$, $R^{\prime}_{k}:=\varphi_{i_{1},j^{\prime}}\cdots\varphi_{i_{k-1},j^{\prime}}\varphi_{i^{\prime}_{k},j^{\prime}}([0,1]^{2})$ is another level-$k$ rectangle such that $\pi_{2}(R^{\prime}_{k})=\pi_{2}(R_{k})$ (i.e, they are on the same row of the $n^{-k}\times m^{-k}$ grid) while $\pi_{1}(R^{\prime}_{k})$ is a $\frac{|i^{\prime}_{k}-i_{k}|}{n^{k}}$-translation of $\pi_{1}(R_{k})$. Note that $\varphi_{Q_{k}}^{-1}f(R^{\prime}_{k})$ is simply a translated copy of $\varphi_{Q_{k}}^{-1}f(R_{k})$ along the line supporting $\varphi_{Q_{k}}^{-1}f(\ell^{k})$, and they have disjoint interiors. Picking $k$ large at the beginning (as long as $|\pi_{1}(\widetilde{E}_{2})|=|\pi_{1}(\widetilde{E}^{\prime}_{2})|=1$), the set
$\widetilde{E}^{\prime}_{2}:=\varphi_{Q_{k}}^{-1}f(R^{\prime}_{k})\cap\{x\in\mathbb{R}^{2}:\pi_{1}(x)\in[-2,-1]\},$
is simply a translated copy of $\widetilde{E}_{2}$ (see Figure 4 for an illustration). By (3.6), the segment $[-2,-1]\times\{\frac{j_{*}}{m}\}$ lies entirely in $\widetilde{E}_{2}$ and hence does not meet the interior of $\widetilde{E}^{\prime}_{2}$. Moreover, by the definition of $R^{\prime}_{k}$, it is not hard to check that either (when $\widetilde{E}^{\prime}_{2}$ lies above that segment)
$\frac{j_{*}}{m}\leq\min\pi_{2}(\widetilde{E}^{\prime}_{2})<\max\pi_{2}(\widetilde{E}^{\prime}_{2})\leq\min\pi_{2}(\widetilde{E}_{2})+2n\lambda<\frac{j_{*}+1}{m};$
or (when $\widetilde{E}^{\prime}_{2}$ lies below that segment)
$\frac{j_{*}}{m}\geq\max\pi_{2}(\widetilde{E}^{\prime}_{2})>\min\pi_{2}(\widetilde{E}^{\prime}_{2})\geq\frac{j_{*}}{m}-2n\lambda>\frac{j_{*}-1}{m}.$
Without loss of generality, let us assume the former. So Condition A holds for $\widetilde{E}^{\prime}_{2}$ and $j_{*}$. Applying the previous argument (i.e., the one after assuming Condition A), one can show that the parallelogram $\varphi_{Q_{k}}^{-1}f(R^{\prime}_{k})$ passes through a hole in $\bigcup_{z\in\mathbb{Z}^{2}}K+z$, which leads to a contradiction as before.
[…]
Figure 4. An illustration of the possible positions of $\widetilde{E}^{\prime}_{2}$ (color online)
Finally, if #$\#I_{j}=1$ for all $j$, then since $n>m\geq 2$, we can find $10\leq i_{*}\leq n-$ 1 such that $(i_{*},j_{*}\pmod{m}),(i_{*},j_{*}-1)\notin\Lambda$. Write $j^{\prime}_{*}:=j_{*}\pmod{m}$. Then, since $\pi_{2}(\widetilde{E}_{2})\subset[\tfrac{j_{*}-1}{m},\tfrac{j_{*}+1}{m}]$, applying the previous argument (i.e., the one after assuming Condition A) to $t_{0}=2$, $j_{0}=j_{*}$, $i_{0}=i_{*}$, $j^{\prime}_{0}=j^{\prime}_{*}$ and $U=(\frac{i_{0}}{n},\frac{i_{0}+1}{n})\times(\frac{j^{\prime}_{*}-1}{m},\frac{j^{\prime}_{*}+1}{m})$, one can show that the parallelogram $\varphi_{Q_{k}}^{-1}f(R_{k})$ passes through a hole (which is a translatied copy of $U$) in $\bigcup_{z\in\mathbb{Z}^{2}}K+z$, which leads to a contradiction as before. ∎